OpenStudy (anonymous):

Find the direct variation equation of the graph through the points (0, 0) and (1, -2). Write in y=kx form. can someone explain this to me?

OpenStudy (anonymous):

please

OpenStudy (anonymous):

let y=mx+c put (0,0), 0=0+c c=0 therefore, y=mx m=slope = (-2-0)/(1-0) = -2 the required equation is y=-2x